OpenStudy (anonymous):

let f(x,y)=(xy^2)/(x^2+y^2), evaluate 0f/0y at the point (1,-2)

zepdrix (zepdrix):

So we need to find the partial derivative of f with respect to y? Is that what the zero's are suppose to be? :)

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\large f(x,y)=\frac{xy^2}{x^2+y^2}\] Hmm so it looks like we need to apply the quotient rule.\[\large \frac{\partial f}{\partial y}=\frac{\color{royalblue}{(xy^2)'}(x^2+y^2)-(xy^2)\color{royalblue}{(x^2+y^2)'}}{(x^2+y^2)^2}\] Understand the setup? We'll need to take the partial derivatives of the blue terms with respect to y.

OpenStudy (anonymous):

Okay how do i do that

zepdrix (zepdrix):

Umm ok let's look at the first one.\[\large \frac{\partial}{\partial y}(xy^2)\]When we take a partial, we hold all other variables constant. So you could think of the problem this way,\[\large \frac{\partial}{\partial y}(cy^2)\]From here we would apply the derivative as normal, and the constant doesn't affect the process.\[\large 2cy\] But our answer is actually, \(\large 2xy\), the c was just to hopefully prove a point :)

zepdrix (zepdrix):

\[\large \frac{\partial f}{\partial y}=\frac{\color{orangered}{(2xy)}(x^2+y^2)-(xy^2)\color{royalblue}{(x^2+y^2)'}}{(x^2+y^2)^2}\] Can you do the other term using this same idea? \[\large \frac{\partial}{\partial y}(c^2+y^2)\]

OpenStudy (anonymous):

Would it be 2y

zepdrix (zepdrix):

So the derivative of the constant became zero? Yes good job :) \[\large f_y(x,y)=\frac{\color{orangered}{(2xy)}(x^2+y^2)-(xy^2)\color{orangered}{(2y)}}{(x^2+y^2)^2}\]

OpenStudy (anonymous):

is that the final answer?

zepdrix (zepdrix):

Don't do any simplification just yet. From here, plug in your coordinate pair that you were given. It should be easier to work out if you plug the numbers before simplifying. So we want this, \(\large f_y(1,-2)\) So plug in 1's for all your x's, and -2's for all your y's.

OpenStudy (anonymous):

-4/25?

zepdrix (zepdrix):

Yay good job! \c:/ That should be our final answer.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Could you also show me how to do (x^2+xy+y)^5

zepdrix (zepdrix):

With respect to y again?

OpenStudy (anonymous):

yes just like the first one

OpenStudy (anonymous):

I got -5

zepdrix (zepdrix):

Oh same coordinates as the first problem? (1,-2)

OpenStudy (anonymous):

(-1,1)

zepdrix (zepdrix):

Hmm I'm getting zero. Lemme double check my work a sec.

OpenStudy (anonymous):

what was the partial dx

zepdrix (zepdrix):

\[\large (x^2+xy+y)^5\] Partial with respect to y,\[\large 5(x^2+xy+y)^4(x+1)\]Understand where the (x+1) is coming from? Partial with respect to x,\[\large 5(x^2+xy+y)^4(2x+y)\]

OpenStudy (anonymous):

Okay that is what i got

zepdrix (zepdrix):

So for the partial y, simply plugging in x=-1 shows us that it is zero.\[\large 5(x^2+xy+y)^4\color{orangered}{(-1+1)}\]

zepdrix (zepdrix):

Oh but for the partial x, yes I'm getting -5 also. I think that's what you were saying earlier. :)

OpenStudy (anonymous):

DO you know how find sketch the level curve f(x,y)=C where f(x,y)=-e^x+y and C=2